# Enhanced Collisional Losses from a Magnetic Mirror Using the Lenard-Bernstein Collision Operator

### v2

@@ 2 Method of Images Solution
where $w=\exp(a^{2}-\bar{v}_{0}^{2})$. This can be inverted to give $a^{2}=\bar{v}^{2}_{0}+\ln(w)$, which determines the edge of the sink region, $a$, in terms of the tip of the loss region $v_{0}$. To calculate this system’s confinement time and energy loss rate, we integrate the image sinks over all velocity space, considering the symmetry in gyro- and pitch-angle.
@@
$\frac{1}{\tau_{c}}=\frac{1}{n_{s}}\frac{dn_{s}}{dt}$ $=-\nu_{sLBO}2Z_{s}\frac{\mathrm{Erfc}(a)}{\ln\left(\frac{w+1}{w-1}\right)},$ (24)
$\frac{1}{E_{s}}\frac{dE_{s}}{dt}=\frac{1}{n_{s}T_{s}}\frac{d(n_{s}T_{s})}{dt}$ $=-\nu_{sLBO}\frac{4}{3\sqrt{\pi}}Z_{s}\frac{e^{-a^{2}}a+\frac{\sqrt{\pi}}{2}\mathrm{Erfc}(a)}{\ln\left(\frac{w+1}{w-1}\right)}.$ (25)
@@ 3 Numerical simulations and corrections to equation (20)
In figure 3, the numerical code is validated by comparing the analytic results presented in Pastukhov (1974) and Najmabadi et al. (1984) with corrections noticed by Cohen et al. (1978), Najmabadi et al. (1984), equation (24), and equation (26), which in turn were validated using other numerical methods. Results in figure 3 agree well with those in many prior works (Najmabadi et al., 1984; Cohen et al., 1978; Fyfe et al., 1981). The confinement time $\tau_{c}$ is normalized to collision frequency for generality. The diamonds are the finite element method numerical results, and the dashed lines are the analytic approximations. Although figure 3 shows great agreement with prior work, it has a poor agreement with equation (24). Najmabadi et al. (1984) overcome these errors by finding a correction coefficient in their equation (41b) based on the difference in flux from the approximate loss cone from the solution $g(\bar{v},\mu)$ versus the real loss cone. We have not replicated the calculation used to determine their equation (41b), so instead, we opt to calculate this correction numerically. We look for a correction that adds a coefficient to equation (20) in the same place that Najmabadi et al. (1984) finds. We call this correction coefficient $c_{0}$ and it is chosen to minimize error with the finite element code. Equations (20), (24), and (25) are modified to be the following.
@@
Table 1: A table of optimal correction factors $c_{0}$ for various values of $z_{s}e\phi/T_{s}$ and $R$ to minimize error with the finite element code for pure electrons with $Z_{s}=1$.
@@
Figure 5(c) shows that our equation (26) and the numerical approach exhibit similar trends, but are very different in magnitude. It is important to note that figure 5(c) subtracts the linear component $z_{s}e\phi/T_{s}$ to highlight the differences between the different results. Comparing results to prior work, our numerical method produces roughly a 20% difference with respect to the analytic results from Najmabadi et al. (1984). Although errors of 20% are large, Najmabadi et al. (1984) noticed similar errors, so these results are comparable to prior work. Interestingly, while the analytic results of Najmabadi et al. (1984) seem to have a constant 20% error, the discrepancy between (26) and the numerical results grows with $z_{s}e\phi/T_{s}$. Similar errors of 20% are observed at low values of $z_{s}e\phi/T_{s}$; at large values, the error grows to nearly 100% at $z_{s}e\phi/T_{s}=8$. The LBO model has an overall higher average loss energy than the model used by Najmabadi et al. (1984). This means the losses from LBO collisions are more spread around the loss hyperboloid. In contrast, a more accurate collision operator would have losses more concentrated around the tip of the loss hyperboloid. It is worth noting that the correction factor $c_{0}$ does not modify equation (26).
@@ 4 Discussion
The scrape-off layer of toroidal confinement devices, like tokamaks and stellarators, can exhibit a Pastukhov potential due to the difference in magnetic field strength between the inboard and outboard sides (Majeski et al., 2017). The magnetic field magnitude ($|B|$) varies inversely with the distance ($R$) from the central axis, determining the mirror ratio ($R_{\mathrm{out}}/R_{\mathrm{in}}$) of these devices, typically around 2. Additionally, observed potentials are approximately $z_{s}e\phi/T\sim 2$. This implies these devices have low mirror ratios and a limited potential. In this scenario, figure 5(a) aligns reasonably with previous findings by Najmabadi et al. (1984). However, Pastukhov (1974) and Najmabadi et al. (1984) consider a high mirror ratio and potential limit, making their results inaccurate for toroidal confinement devices. Caution is necessary when applying these equations in regimes approaching the bounds of these limits. The variational method for studying collisional losses not considered here offers an approach suitable for arbitrary mirror ratios and ambipolar potential, making results from Catto & Li (1985) and Khudik (1997) more appropriate.
We note a subtlety in the problem setup as defined by Pastukhov, and Najmabadi, and others. It is usually stated that there is a source of particles and energy at low energy to balance the sink at high energy. Standard electron-electron collision operators are energy conserving, and the electron-ion collision operator has been approximated by pitch angle scattering, which also conserves energy, so one may wonder where the energy injection comes from. It should be noted that it is not sufficient to add a source to the kinetic equation of the form $S(v)=S_{0}\delta(v-v_{I})/(4\pi v^{2})$ (where $S_{0}$ is the source rate in units of particles per second per unit volume), because no choice of the injection velocity $v_{I}$ can provide enough power for steady state unless $v_{I}$ is quite large, violating the assumption that the source is at low energy. This is because the source energy per injected particle, $(1/2)mv_{I}^{2}$, must be the same in steady state as the average energy per particle lost, which from (27) is $E_{\rm loss}\sim z_{s}e\phi\gg T_{e}$.
@@ Appendix A Convenient choice for the free function $q(\bar{v})$
In Najmabadi et al. (1984), they determine $c_{N}=0.84$, although we have not replicated the calculation used to determine this. In equation (37), it is essential to notice that the correction coefficient does change the asymptotic convergence of the problem. The authors claim that this convergence is enhanced from $\mathcal{O}(1/x_{a}^{2})$ to $\mathcal{O}(\exp(-x_{z}^{2}))$, but adding this 0.84 changes where $q_{0}$ asymptotes to, so it does more than yield faster convergence; it changes the convergence entirely. This is showcased in figure 6, where it is clear that the error asymptotes to a constant level, which depends strongly on $c_{N}$. Without using this correction factor, the error saturates to around $40\%$ compared to the numerical code, which is quite large and not converging. With $c_{N}=0.84$, the finite element code has roughly a 10% error with equation 37, whereas Najmabadi et al. (1984) reports errors of ±$\pm 7\%$. Their expression for $q_{0}$ is the same as with the LBO, $\ln(w+1/w-1)$, but their $w$ takes the form for large $z_{s}e\phi/T_{s}$ of $w=\sqrt{1+1/RZ_{s}}$. Therefore $q_{0}$ asymptotes at large R to $\ln(1+4RZ_{s})$. This number is not large and is comparable to the coefficient of $0.84$ they find.
@@
Table 2: A table of optimal correction factors $c_{N}$ for various values of $z_{s}e\phi/T_{s}$ and $R$ to minimize error with the finite element code for ions $Z_{s}=1/2$.
@@
Table 3: A table of optimal correction factors $c_{N}$ for various values of $z_{s}e\phi/T_{s}$ and $R$ to minimize error with the finite element code for pure electrons $Z_{s}=1$.

### v3

@@ 2 Method of Images Solution
where $w=\exp(a^{2}-\bar{v}_{0}^{2})$. This can be inverted to give $a^{2}=\bar{v}^{2}_{0}+\ln(w)$, which determines the edge of the sink region, $a$, in terms of the tip of the loss region $v_{0}$.
In Najmabadi et al. (1984), they adjust the strength of the sinks by examining the difference between the flux of the true loss cone, equation (1), and the approximate loss cone, where equation (19) equals zero. They account for this in equations 41a and 41b. Their instructions are to evaluate (19) along the loss cone one $z_{s}e\phi/T_{s}$ above the tip of the loss cone in the limit where $R\gg 1$. This means $\bar{v}^{2}=v_{0}^{2}+1$, $\mu=1$, and $a\approx v_{0}$, which leads to $z=\exp(v_{0}^{2}+1)$ and $\rho=0$, meaning $g(0,v_{0}^{2}+1)\approx 1-q_{0}\ln((e+1)/(e-1))=1-0.77q_{0}$. Thus, we modify $q_{0}$ by subtracting $0.77$, leading to a corrected definition of $q_{0}$. Najmabadi et al. (1984) determines this to be $0.84q_{0}$, but we have not replicated the calculation used to determine their equation (41b).
$q_{0}$ $=\left(\ln\left(\frac{w+1}{w-1}\right)-0.77\right)^{-1}.$ (22)
To calculate this system’s confinement time and energy loss rate, we integrate the image sinks over all velocity space, considering the symmetry in gyro- and pitch-angle.
@@
$\frac{1}{\tau_{c}}=\frac{1}{n_{s}}\frac{dn_{s}}{dt}$ $=-\nu_{sLBO}2Z_{s}\frac{\mathrm{Erfc}(a)}{\ln\left(\frac{w+1}{w-1}\right)-0.77},$ (25)
$\frac{1}{E_{s}}\frac{dE_{s}}{dt}=\frac{1}{n_{s}T_{s}}\frac{d(n_{s}T_{s})}{dt}$ $=-\nu_{sLBO}\frac{4}{3\sqrt{\pi}}Z_{s}\frac{e^{-a^{2}}a+\frac{\sqrt{\pi}}{2}\mathrm{Erfc}(a)}{\ln\left(\frac{w+1}{w-1}\right)-0.77}.$ (26)
@@ 3 Numerical simulations and corrections to equation (20)
In figure 3, the numerical code is validated by comparing the analytic results presented in Pastukhov (1974) and Najmabadi et al. (1984) with corrections noticed by Cohen et al. (1978), Najmabadi et al. (1984), equation (25), and equation (27), which in turn were validated using other numerical methods. Results in figure 3 agree well with those in many prior works (Najmabadi et al., 1984; Cohen et al., 1978; Fyfe et al., 1981). The confinement time $\tau_{c}$ is normalized to collision frequency for generality. The diamonds are the finite element method numerical results, and the dashed lines are the analytic approximations. Although figure 3 shows great agreement with prior work, it has a fair agreement with equation (25). To improve agreement between the code and equation (25), rather than calculating the flux correction coefficient by hand to be $0.77$ in (22), we can calculate this number numerically. First, let’s call this correction coefficient $c_{0}$. It can be chosen to minimize error with the finite element code. Equations (20), (25), and (26) are modified to be the following.
@@
Table 1: A table of optimal correction factors $c_{0}$ for various values of $z_{s}e\phi/T_{s}$ and $R$ to minimize error with the finite element code with $Z_{s}=1$.
@@
Figure 5(c) shows that our equation (27) and the numerical approach exhibit similar trends, but are very different in magnitude. It is important to note that figure 5(c) subtracts the linear component $z_{s}e\phi/T_{s}$ to highlight the differences between the different results. Comparing results to prior work, our numerical method produces roughly a 20% difference with respect to the analytic results from Najmabadi et al. (1984), investigated in appendix B. Although errors of 20% are large, Najmabadi et al. (1984) noticed similar errors, so these results are comparable to prior work. Interestingly, while the analytic results of Najmabadi et al. (1984) seem to have a constant 20% error, the discrepancy between (27) and the numerical results grows with $z_{s}e\phi/T_{s}$. With the LBO at $c_{0}=0.77$, similar errors of 20% are observed at low values of $z_{s}e\phi/T_{s}$; at large values, the error grows to nearly 50% at $z_{s}e\phi/T_{s}=8$. The LBO model has an overall higher average loss energy than the model used by Najmabadi et al. (1984). This means the losses from LBO collisions are more spread around the loss hyperboloid. In contrast, a more accurate collision operator would have losses more concentrated around the tip of the loss hyperboloid. It is worth noting that the correction factor $c_{0}$ does not modify equation (27).
@@ 4 Discussion
The scrape-off layer of toroidal confinement devices, like tokamaks and stellarators, can exhibit a Pastukhov potential due to the difference in magnetic field strength between the inboard and outboard sides (Majeski et al., 2017). The magnetic field magnitude ($|B|$) varies inversely with the distance ($R$) from the central axis, determining the mirror ratio ($R_{\mathrm{out}}/R_{\mathrm{in}}$) of these devices, typically around 2. Additionally, observed potentials are approximately $z_{s}e\phi/T\sim 2$. This implies these devices have low mirror ratios and a limited potential. In this scenario, figure 5(a) aligns reasonably with previous findings by Najmabadi et al. (1984). However, Pastukhov (1974) and Najmabadi et al. (1984) consider a high mirror ratio and potential limit, making their results inappropriate for toroidal confinement devices. Caution is necessary when applying these equations in regimes approaching the bounds of these limits. The variational method for studying collisional losses not considered here offers an approach suitable for arbitrary mirror ratios and ambipolar potential, making results from Catto & Li (1985) and Khudik (1997) more appropriate.
We note a subtlety in the problem setup as defined by Pastukhov, Najmabadi, and others. It is usually stated that there is a low-energy source of particles to balance the sink at high energy. Standard electron-electron collision operators are energy conserving, and the electron-ion collision operator has been approximated by pitch angle scattering, which also conserves energy, so one may wonder where the energy injection comes from. It should be noted that it is not sufficient to add a source to the kinetic equation of the form $S(v)=S_{0}\delta(v-v_{I})/(4\pi v^{2})$ (where $S_{0}$ is the source rate in units of particles per second per unit volume), because no choice of the injection velocity $v_{I}$ can provide enough power for steady state unless $v_{I}$ is quite large, violating the assumption that the source is at low energy. This is because the source energy per injected particle, $(1/2)mv_{I}^{2}$, must be the same in steady state as the average energy per particle lost, which from (28) is $E_{\rm loss}\sim z_{s}e\phi\gg T_{e}$.
@@ Appendix A Convenient choice for the free function $q(\bar{v})$
In Najmabadi et al. (1984), they determine $c_{N}=0.84$, although we have not replicated the calculation used to determine this. In equation (38), it is essential to notice that the correction coefficient does change the asymptotic convergence of the problem. The authors claim that this convergence is enhanced from $\mathcal{O}(1/x_{a}^{2})$ to $\mathcal{O}(\exp(-x_{z}^{2}))$, but adding this 0.84 changes where $q_{0}$ asymptotes to, so it does more than yield faster convergence; it changes the convergence entirely. This is showcased in figure 6, where it is clear that the error asymptotes to a constant level, which depends strongly on $c_{N}$. Without using this correction factor, the error saturates to around $40\%$ compared to the numerical code, which is quite large and not converging. With $c_{N}=0.84$, equation 38 has a roughly 10% error with the finite element code, whereas Najmabadi et al. (1984) reports errors of ±$\pm 7\%$. Their expression for $q_{0}$ is the same as with the LBO, $\ln(w+1/w-1)$, but their $w$ takes the form for large $z_{s}e\phi/T_{s}$ of $w=\sqrt{1+1/RZ_{s}}$. Therefore $q_{0}$ asymptotes at large R to $\ln(1+4RZ_{s})$. This number is not large and is comparable to the coefficient of $0.84$ they find.
@@
Table 2: A table of optimal correction factors $c_{N}$ for various values of $z_{s}e\phi/T_{s}$ and $R$ to minimize error with the finite element code for hydrogen $Z_{s}=1/2$.
@@
Table 3: A table of optimal correction factors $c_{N}$ for various values of $z_{s}e\phi/T_{s}$ and $R$ to minimize error with the finite element code for electrons $Z_{s}=1$.